# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v1

@@ Abstract
In this paper, we aim to discuss three key questions:
  • Q1. Justification of the model of relaxation time approximation (RTA) mathematically;
  • Q2. Pole/cut structure in the retarded correlators within the kinetic description;
  • Q3. Using only physical arguments and the eigenspectrum information of the linearized
    Boltzmann collision operator $\mathcal{L}_{0}$ defined in the main text to obtain the eigenspectrum
    properties of a class of redefined linearized collision operators.
Some are interrelated, and all hinge on the properties of the eigenspectrum of $\mathcal{L}_{0}$. Understanding whether a specific interacting system permits the RTA directly decides the pole/cut structure in the retarded correlation function, revealing the equivalence between Q1 and Q2.
In answering Q1, we find that only the RTA with an energy-independent relaxation time can be justified in the case of hard interactions. Consequently, the quest for a well-defined relaxation time approximation involves testing various interactions to identify the rare hard cases, such as a gas of elastically colliding neutrinos interacting through a four-fermion vertex. We also propose an alternative approach to restore the collision invariance lacking in traditional RTA by truncating the full linearized collision operator, rather than introducing counter terms.
Regarding Q2, our findings in the vanishing mass limit are as follows. For hard interactions, the eigenspectrum of the linearized collision operator is gapped, with hydrodynamic poles below this gap being the dominating non-analytical structures. Whereas for soft interactions, commonly encountered in relativistic kinetic theory, the eigenspectrum becomes gapless, leading to branch-cuts emerging as the dominant non-analytical structures
$-\infty<\text{Im }\omega<0,\quad-k\leq\text{Re}\,\omega\leq k,$
consistent with existing studies in the vanishing $k$ limit. When particles constituting the system are massive, the non-analytical structures become more complex and richer.
When it comes to Q3, we demonstrate that physical arguments and knowledge of $\mathcal{L}_{0}$’s eigenspectrum can guide us in seeking the eigenspectrum properties of a class of redefined linearized collision operators. From our perspective, given the extreme challenge of mathematically analyzing the redefined linearized Boltzmann integral operator, the ability to rule it with intuitive arguments is highly nontrivial.
@@ I Introduction
Over recent decades, relativistic hydrodynamics has notably contributed to characterizing the dynamic evolution of Quark-Gluon Plasma (QGP), as observed in heavy-ion collision experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) [7]. Within the framework of phenomenological modeling for these collisions, the initial temporal parameter, crucial for initiating hydrodynamics, has been meticulously minimized. This adjustment, made in accordance with experimental observations, suggests that the system may rapidly equilibrate. Moreover, the applicability of hydrodynamics has been noted to extend into less expected domains, including smaller collision systems like nucleus-nucleon and proton-proton collisions, as evidenced by recent studies [8, 9]. What is the dynamic mechanism that triggers rapid equilibration? Why does relativistic hydrodynamics also work well for small systems? Regarding the first question, it has been proposed that hydrodynamics can be viewed as an attractor governing the late-time behavior of systems as they approach equilibrium [10]. At the early stage of evolution, the system tends to flow towards the hydrodynamic attractor, even when far from local equilibrium, which may account for rapid hydrodynamization. Thus, hydrodynamization may have a broader application range than thermalization, contrasting with the conventional view that hydrodynamics is a truncated gradient expansion near local equilibrium. Despite significant advancements, it remains an open question how hydrodynamization evolves from a general dynamic system with diverse microscopic interactions.
@@
As the first step towards understanding how relativistic nonequilibrium systems reach thermal equilibrium, the properties of retarded correlation functions have recently garnered extensive research interests. Two-point retarded correlation functions are pivotal, as they provide a wealth of insights into the transport characteristics of multi-particle systems, particularly how an equilibrium system responds to off-equilibrium disturbances within the linear regime. Moreover, non-analytical aspects—such as poles or discontinuities in Fourier space—are crucial for determining the system’s evolutionary patterns. Poles are indicative of collective excitations that evolve towards equilibrium, corresponding to the persistent hydrodynamic modes, whereas the presence of cuts or non-hydrodynamic modes is intrinsically linked to the emergence of hydrodynamic phenomena and the applicability of hydrodynamics. Research has shown that the correlators contain only poles at infinite t’Hooft coupling in large $N$ thermal $\mathcal{N}=4$ Super Yang-Mills (SYM) theory. In the holographic description, the spectra of correlation functions correspond to the ring-down spectra of dual linearly perturbed black holes: the quasinormal modes [11, 12]. However, in Ref. [11], the authors argue that the branch-cut structure emerges in the regime of weak, but finite t’Hooft coupling, indicating a possible transition behavior controlled by the t’Hooft coupling. Motivated by this illuminating finding, Paul Romatschke initiated research on thermal correlators of large $N$ gauge theories in effective kinetic theory [1](see also [13]), reporting the onset of transition behavior for hydrodynamic poles. In subsequent related studies, including analytical estimates [14], qualitative models [15], and numerical calculations [2, 6], the dominant non-analytical structure is found to be the branch-cut rather than poles, thus posing the ’poles or cuts?’ dilemma. This is one of the topics we aim to explore in this manuscript, corresponding to Q2 as introduced in the abstract.
@@
As a universal low-energy effective theory, hydrodynamics describes the collective, macroscopic dynamics over large distances and time scales. As previously mentioned, elucidating the relationship between the microscopic dynamics of matter’s fundamental constituents and the macroscopic dynamics of coarse-grained degrees of freedom is a fundamental question. Typically, collective macroscopic dynamics heavily relies on the intricate microscopic dynamic details involving a vast number of degrees of freedom: for instance, a microscopic theory is required to determine the relevant coefficients in the macroscopic description. In principle, an accurate description of a dynamic system necessitates a comprehensive treatment of all particles at the microscopic level. However, tracking the evolution of all particles is impractical; thus, a reduced description of microscopic degrees of freedom is essential. In scenarios of weak coupling, where the concept of quasiparticles is applicable, kinetic theory serves as an effective tool for describing many-body systems, with the relevant degrees of freedom being statistical distribution functions. For example, the BBGKY hierarchy equation describes the coupled dynamic evolution of n-particle distribution functions [16, 17, 18, 19] . The BBGKY hierarchy equation is an infinite tower of integro-differential equations, making it challenging to solve without some form of truncation. By disregarding n-particle correlations (for $n\geq 2$) and truncating this infinite series of integro-differential equations to the lowest order, we recover the renowned Boltzmann equation, which provides the evolution of the one-particle distribution function. Despite its omission of nearly all particle correlations, the Boltzmann equation maintains profound physical significance and is capable of characterizing various non-equilibrium phenomena within kinetic theory. Notably, hydrodynamic behavior can be derived from the long-wavelength limit of the Boltzmann equation. Hence, the Boltzmann equation is an excellent candidate, particularly for weakly coupled cases, for elucidating how hydrodynamization is achieved in a dynamic system and how the multitude of microscopic degrees of freedom condense into coarse-grained hydrodynamic degrees of freedom.
@@
However, the Boltzmann equation contains an intractable collision integral, rendering its analytical or numerical solution challenging. Even with the simplest interactions, such as the hard-sphere potential, the linearized Boltzmann equation’s collision operator retains a complex structure, making its analysis extremely difficult. Recently, the eigenspectrum of the linearized Boltzmann collision operator in massless scalar $\phi^{4}$ theory was analytically determined [5]. However, extending this to more realistic interactions remains a significant challenge [^1] The first analytical eigenspectrum was obtained by C.S.Wang Chang and U.E.Uhlenbeck in the context of monatomic gases in the non-relativistic case (see chapter IV of [20]).. In 1974, Anderson and Witting (AW) proposed an approximation of the relativistic Boltzmann equation using a simplified collision operator [21], effectively extending the BGK (Bhatnagar, Gross, and Krook) model to the relativistic domain [22]. The AW model, also known as the relaxation time approximation, omits much of the dynamic information found in the full collision operator. Firstly, it disregards nonlinearity. Secondly, its validity hinges on a clear separation between the eigenvalue representing the slowest relaxation and the others. Despite these limitations, the relaxation time approximation addresses practical challenges, enabling analytic and semi-analytic solutions for the simplified Boltzmann equation. This approach paves the way for an insightful and instructive analysis. Furthermore, the AW RTA model has been extended to a more general form with nontrivial energy dependence. However, the justification of this extended RTA model within kinetic theory awaits rigorous validation, which is another topic of interest, referred to as Q1 in the abstract.
@@ II Linearized Boltzmann equation
As mentioned before, the complicated Boltzmann equation is often linearized to facilitate a more straightforward analysis. Following the linearization procedure, we expand the distribution function around the local equilibrium state $f(x,p)=f_{0}(x,p)(1+\chi(x,p)\,)$. Consequently, Eq.(1) transforms into the following form
$p\cdot\partial f(x,p)=-f_{0}(x,p)\mathcal{L}_{1}[\chi],$ (3)
@@
$-\mathcal{L}_{1}[\chi]\equiv\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{0}(x,p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}\big{(}\chi(x,p_{1})+\chi(x,p_{2})-\chi(x,p)-\chi(x,p^{\prime})\big{)}.$ (4)
The local equilibrium distribution is defined as
@@
where $\beta^{\mu}\equiv\frac{u^{\mu}}{T},\xi\equiv\frac{\mu}{T},\beta\equiv\frac{1}{T}$ with the local temperature $T(x)$, and the chemical potential $\mu(x)$ associated with the conserved particle number. By construction, leveraging the collisional invariance of 1 and $p_{\mu}$, the local equilibrium distribution nullifies the collision kernel on the right-hand side (RHS) of Eqs.(1) and (3): $C[f_{0}]=\mathcal{L}_{1}[f_{0}]=0$. From a mathematical standpoint, it can be verified that the linearized collision operator $\mathcal{L}_{1}$ is self-adjoint and positive semidefinite within the square-integrable Hilbert space [24],
$\int{\rm dP}f_{0}(p)\psi(p)\mathcal{L}_{1}\phi(p)=\int{\rm dP}f_{0}(p)\phi(p)\mathcal{L}_{1}\psi(p),\quad\int{\rm dP}f_{0}(p)\psi(p)\mathcal{L}_{1}\psi(p)\geq 0,$ (6)
@@
In certain special cases, an additional simplification proves particularly useful and pertinent. It should be noted that the local equilibrium distribution does not constitute a solution to the Boltzmann equation. However, the global equilibrium distribution function $f_{eq}(p)$, a specific instance of a local equilibrium distribution, does solve the Boltzmann equation. When it comes to the normal mode solution of the Boltzmann equation, the background around which the distribution function is expanded is typically assumed to be a homogeneous and static equilibrium configuration [1, 26]. For the sake of subsequent discussions, we introduce this simplification as well. Adopting the expansion $f=f_{eq}(1+\chi(x,p)\,)$, Eq.(3) can be further simplified to
$\partial_{t}\chi(x,p)+\bm{v}\cdot\nabla\chi(x,p)=-p_{0}^{-1}\mathcal{L}_{1}[\chi],$ (7)
where $\bm{v}\equiv\frac{\bm{p}}{p_{0}}$ and we should change $f_{0}(x,p)$ into $f_{eq}(p)$ accordingly.
Also we can restate Eq.(3) as
where $D\equiv u\cdot\partial$, $E_{p}=u\cdot p$ and $p^{\langle\nu\rangle}\equiv\Delta^{\nu\rho}p_{\rho}$. Here we introduce a new collision operator $\mathcal{L}_{0}\equiv E_{p}^{-1}\mathcal{L}_{1}$, which retains the properties of being positive semidefinite and self-adjoint
$\int{\rm dP}f_{0}(p)E_{p}\psi(p)\mathcal{L}_{0}\phi(p)=\int{\rm dP}f_{0}(p)E_{p}\phi(p)\mathcal{L}_{0}\psi(p),\quad\int{\rm dP}f_{0}(p)E_{p}\psi(p)\mathcal{L}_{0}\psi(p)\geq 0.$ (9)
Note the weight function transforms into $f_{0}(p)E_{p}$. Similarly, we have
@@
as the counterpart of Eq.(7).
@@ III Revisiting the relaxation time approximation
where $\mathcal{L}_{\alpha}\equiv E_{p}^{\alpha}\mathcal{L}_{0}$ represents the redefined linearized collision operator with $\mathcal{L}_{1}$ and $\mathcal{L}_{0}$ being specific instances. It is easily proved that $\mathcal{L}_{\alpha}$ inherits the positive semidefinite and self-adjoint properties from $\mathcal{L}_{0}$. This can be achieved by redefining the weight function within the inner product definition to be $f_{0}(x,p)E_{p}^{1-\alpha}$ in Eq.(9). In the following subsections III.2, III.2 and III.3, we illustrate that naively truncating $\mathcal{L}_{\alpha}$ leads to a form resembling the energy-dependent RTA; however the resulting RTA-like model is not well-defined.
@@ III.1 Anderson and Witting model
which can be formally solved to get $\chi(t)=e^{-\mathcal{L}_{0}t}\chi(0)$. Since $\mathcal{L}_{0}$ is self-adjoint and semipositive, the linear perturbation from equilibrium decays exponentially, which is consistent with stability requirements. The same working logic applied to Eq.(14) leads to $\chi^{\prime}(t)=e^{-t/\tau_{R}}\chi^{\prime}(0)$. By comparison, it is natural to interpret $\tau_{R}$ as the relaxation time for the linear perturbation from equilibrium, and $\mathcal{L}_{0}^{-1}$ as the infinite-dimensional relaxation time matrix (we adopt the terminology ”matrix” although it may not be appropriate to use it when the eigenvalue spectrum is continuous).
@@
For comparative purposes, we also present the expansion form of $\chi^{\prime}=\sum_{n}c^{\prime}_{n}(t)\psi_{n}(p)$. Additionally, the counter term $-\sum_{n=1}^{5}c^{\prime}_{n}(0)\psi_{n}(p)e^{-t/\tau_{R}}$ is introduced to account for the absence of collision invariants in the traditional RTA, as further discussed in Sec.V. Note $C(p),C^{\prime}(p)$ are integration constants. In the above equations, the eigenvalues are sequenced in ascending order when increasing $n$. As clearly seen in Eq.(17), the collision invariants, $\psi_{n}(n=1,\cdots 5)$ with zero eigenvalues $\gamma_{n}=0$, contribute to the deviation function $\chi$ but remain unchanged with time. In the solution of Eq.(17), the second term, reflecting the contribution from collision invariants, can also be regarded as an integration constant. By comparing Eqs.(17) and (18), we can deduce that the traditional RTA simplifies the model by condensing the entire nonzero eigenvalue spectrum into a single representative eigenvalue, ${\gamma_{n}}\rightarrow\frac{1}{\tau_{R}}$. Given the exponentially decaying form, $1/\tau_{R}$ should be identified as the smallest eigenvalue, where the mode $\psi_{6}$ persists until the late stage of evolution. In other words, other modes with a larger eigenvalue, which should have been absent from an earlier time, extend their lifetime to the final stage in this approximation.

### v2

@@ Abstract
In this paper, we aim to discuss the following two questions:
  • Q1. Mathematical justification of the model of relaxation time approximation (RTA);
  • Q2. Pole/cut structure in the retarded correlators within the kinetic description.
We find that only the RTA with an energy-independent relaxation time can be justified in the case of hard interactions. Accordingly, we propose an alternative approach to restore the collision invariance lacking in traditional RTA. Furthermore, hydrodynamic poles are the long-lived non-analytical structures in this case. Whereas for soft interactions, commonly encountered in relativistic kinetic theory, the gapless eigenspectrum leads to dominant long-lived branch-cuts. When particles are massive or the perturbations are inhomogeneous, the non-analytical structures become more complex and richer.
@@ I Introduction
Over the past several decades, relativistic hydrodynamics has notably contributed to characterizing the dynamic evolution of Quark-Gluon Plasma (QGP), as observed in heavy-ion collision experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) [1]. Within the framework of phenomenological modeling for these collisions, the initial temporal parameter, crucial for initiating hydrodynamics, has been meticulously minimized to match experimental observations. This adjustment suggests that the system may rapidly equilibrate. Moreover, the applicability of hydrodynamics has been noted to extend into less expected domains, including small collision systems like nucleus-nucleon and proton-proton collisions, as evidenced by recent studies [2, 3]. What is the dynamic mechanism that triggers rapid equilibration? Why does relativistic hydrodynamics also work well for small systems? Regarding the first question, it has been proposed that hydrodynamics can be viewed as an attractor governing the late-time behavior of systems as they approach equilibrium [4]. At the early stage of evolution, the system tends to flow towards the hydrodynamic attractor, even when far from local equilibrium, which may account for rapid hydrodynamization. Thus, hydrodynamization may have a broader application range than thermalization, contrasting with the conventional view that hydrodynamics is a truncated gradient expansion near local equilibrium. Despite significant advancements, it remains an open question how hydrodynamization emerges from a general dynamic system with diverse microscopic interactions.
@@
As the first step towards understanding how relativistic nonequilibrium systems reach thermal equilibrium, the properties of retarded correlation functions have recently garnered extensive research interests. Two-point retarded correlation functions are pivotal, as they provide a wealth of insights into the transport characteristics of multi-particle systems, particularly how an equilibrium system responds to off-equilibrium disturbances within the linear regime. Moreover, non-analytical structures—such as poles or discontinuities in Fourier space—are crucial for determining the system’s evolutionary patterns. Poles are indicative of collective excitations that evolve towards equilibrium, corresponding to the persistent hydrodynamic modes, whereas the presence of cuts or non-hydrodynamic modes is intrinsically linked to the emergence of hydrodynamic phenomena and the applicability of hydrodynamics. Research has shown that the correlators contain only poles at infinite t’Hooft coupling in large $N$ thermal $\mathcal{N}=4$ Super Yang-Mills (SYM) theory. In the holographic description, the spectra of correlation functions correspond to the ring-down spectra of dual linearly perturbed black holes: the quasinormal modes [5, 6]. However, in Ref. [5], the authors argue that the branch-cut structure emerges in the regime of weak, but finite t’Hooft coupling, indicating a possible transition behavior controlled by the t’Hooft coupling. Motivated by this illuminating finding, Paul Romatschke initiated research on thermal correlators of large $N$ gauge theories in effective kinetic theory [7](see also [8]), reporting the onset of transition behavior for hydrodynamic poles. In subsequent related studies, including analytical estimates [9], qualitative models [10, 11], and numerical calculations [12, 13], the dominant non-analytical structure is found to be the branch-cut rather than poles, thus posing the ’poles or cuts?’ dilemma. Specifically, this dilemma initially refers to the mathematical essence of non-hydrodynamic excitations. In this work, we also use it to the denote the comparison of the lifetime of hydrodynamic/non-hydrodynamic excitations. This is one of the topics we aim to explore in this manuscript, corresponding to Q2 as introduced in the abstract.
@@
As a universal low-energy effective theory, hydrodynamics describes the collective, macroscopic dynamics over large distances and time scales. As previously mentioned, elucidating the relationship between the microscopic dynamics of matter’s fundamental constituents and the macroscopic dynamics of coarse-grained degrees of freedom is a fundamental question. Typically, collective macroscopic dynamics heavily relies on the intricate microscopic dynamic details involving a vast number of degrees of freedom: for instance, a microscopic theory is required to determine the relevant coefficients in the macroscopic description. In principle, an accurate description of a dynamic system necessitates a comprehensive treatment of all particles at the microscopic level. However, tracking the evolution of all particles is impractical; thus, a reduced description of microscopic degrees of freedom is essential. In scenarios of weak coupling, where the concept of quasiparticles is applicable, kinetic theory serves as an effective tool for describing many-body systems, with the relevant degrees of freedom being statistical distribution functions. For example, the BBGKY hierarchy equation describes the coupled dynamic evolution of n-particle distribution functions [14, 15, 16, 17] . The BBGKY hierarchy equation is an infinite tower of integro-differential equations, making it challenging to solve without truncation. By disregarding n-particle correlations (for $n\geq 2$) and truncating this infinite series of integro-differential equations to the lowest order, the renowned Boltzmann equation is recovered, which describes the evolution of the one-particle distribution function. Despite its omission of nearly all particle correlations, the Boltzmann equation maintains profound physical significance and is capable of characterizing various non-equilibrium phenomena within kinetic theory. Notably, hydrodynamic behavior can be derived from the long-wavelength limit of the Boltzmann equation. Hence, the Boltzmann equation is an excellent candidate, particularly for weakly coupled cases, for elucidating how hydrodynamization is achieved in a dynamic system and how the multitude of microscopic degrees of freedom condense into coarse-grained hydrodynamic degrees of freedom.
@@
However, the Boltzmann equation still contains an intractable collision integral, rendering its analytical or numerical solution challenging. Even with the simplest interactions, such as the hard-sphere potential, the linearized Boltzmann equation’s collision operator retains a complex structure, making its analysis extremely difficult. Recently, the eigenspectrum of the linearized Boltzmann collision operator in massless scalar $\phi^{4}$ theory was analytically determined [18]. However, extending this to other realistic interactions remains a significant challenge [^1] The first analytical eigenspectrum was obtained by C.S.Wang Chang and U.E.Uhlenbeck in the context of monatomic gases in the non-relativistic case (see chapter IV of [19]).. In 1974, Anderson and Witting (AW) proposed an approximation of the relativistic Boltzmann equation using a simplified collision operator [20], effectively extending the BGK (Bhatnagar, Gross, and Krook) model to the relativistic domain [21]. The AW model, also known as the relaxation time approximation, omits much of the dynamic information in the full collision operator. Firstly, it disregards nonlinearity. Secondly, its validity hinges on a clear separation between the eigenvalue representing the slowest relaxation and the others. Despite these limitations, the relaxation time approximation addresses practical challenges, enabling analytic and semi-analytic solutions for the simplified Boltzmann equation. This approach paves the way for an insightful and instructive analysis. Furthermore, the AW RTA model has been extended to a more general form with nontrivial energy dependence. However, the justification of this extended RTA model within kinetic theory awaits rigorous validation, which is another topic of interest, referred to as Q1 in the abstract.
@@ II Linearized Boltzmann equation
As mentioned before, the complicated Boltzmann equation is often linearized to facilitate a more straightforward analysis. Following the linearization procedure, we expand the distribution function around the local equilibrium state $f(x,p)=f_{0}(x,p)(1+\chi(x,p)\,)$, then Eq.(1) transforms into the following form,
@@
$-\mathcal{L}_{0}[\chi]$ $\equiv E_{p}^{-1}\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{0}(x,p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}$
$\big{(}\chi(x,p_{1})+\chi(x,p_{2})-\chi(x,p)-\chi(x,p^{\prime})\big{)},$ (4)
where $D\equiv u\cdot\partial$, $E_{p}=u\cdot p$ and $p^{\langle\nu\rangle}\equiv\Delta^{\nu\rho}p_{\rho}$. The local equilibrium distribution is defined as
@@
where $\beta^{\mu}\equiv\frac{u^{\mu}}{T},\xi\equiv\frac{\mu}{T},\beta\equiv\frac{1}{T}$ with the local temperature $T(x)$, and the chemical potential $\mu(x)$ associated with the conserved particle number. Evidently, the collisional invariance of 1 and $p_{\mu}$ is respected by construction. From a mathematical standpoint, it can be verified that the linearized collision operator $\mathcal{L}_{0}$ is self-adjoint and positive semidefinite within the square-integrable Hilbert space [23],
$\int{\rm dP}f_{0}(p)E_{p}\psi(p)\mathcal{L}_{0}\phi(p)=\int{\rm dP}f_{0}(p)E_{p}\phi(p)\mathcal{L}_{0}\psi(p),$
$\int{\rm dP}f_{0}(p)E_{p}\psi(p)\mathcal{L}_{0}\psi(p)\geq 0,$ (6)
@@
In certain special cases, an additional simplification proves particularly useful. When it comes to the normal mode solution of the kinetic equation, the background around which the distribution function is expanded is typically assumed to be a homogeneous and static equilibrium configuration $f_{eq}(p)$. For the sake of subsequent discussions, we introduce this simplification as well. Adopting the expansion $f(x,p)=f_{eq}(p)(1+\chi(x,p)\,)$, Eq.(3) can be further simplified to
@@
where $\bm{v}\equiv\frac{\bm{p}}{p_{0}},\nabla_{\alpha}\equiv\Delta^{\beta}_{\alpha}\partial_{\beta}$, and we should change $f_{0}(x,p)$ into $f_{eq}(p)$ in the above expressions accordingly.
@@ III Revisiting the relaxation time approximation
where $\mathcal{L}_{\alpha}\equiv E_{p}^{\alpha}\mathcal{L}_{0}$ represents the redefined linearized collision operator with $\mathcal{L}_{1}$ and $\mathcal{L}_{0}$ being specific instances. It is easily proved that $\mathcal{L}_{\alpha}$ inherits the positive semidefinite and self-adjoint properties from $\mathcal{L}_{0}$. This can be achieved by redefining the weight function within the inner product definition as $f_{0}(x,p)E_{p}^{1-\alpha}$ in Eq.(II). In the following subsections III.2, III.2 and III.3, we illustrate that naively truncating $\mathcal{L}_{\alpha}$ leads to a form resembling the energy-dependent RTA; however the resulting RTA-like model is not well-defined.
@@ III.1 Anderson and Witting model
which can be formally solved to get $\chi(t)=e^{-\mathcal{L}_{0}t}\chi(0)$. Since $\mathcal{L}_{0}$ is self-adjoint and semipositive, the linear perturbation from equilibrium decays exponentially, which is consistent with stability requirements. The same working logic applied to Eq.(11) leads to $\chi^{\prime}(t)=e^{-t/\tau_{R}}\chi^{\prime}(0)$. By comparison, it is natural to interpret $\tau_{R}$ as the relaxation time for the linear perturbation from equilibrium, and $\mathcal{L}_{0}^{-1}$ as the infinite-dimensional relaxation time matrix (we adopt the terminology “matrix” although it may not be appropriate to use it when the eigenvalue spectrum is continuous).
@@
For comparative purposes, we also present the expansion form of $\chi^{\prime}=\sum_{n}c^{\prime}_{n}(t)\psi_{n}(p)$. Additionally, the counter term $-\sum_{n=1}^{5}c^{\prime}_{n}(0)\psi_{n}(p)e^{-t/\tau_{R}}$ is introduced to account for the absence of collision invariance in the traditional RTA, as further discussed in Sec.V. Note $C(p),C^{\prime}(p)$ are integration constants. In the above equations, the eigenvalues are sequenced in ascending order when increasing $n$. As clearly seen in Eq.(14), the collision invariants, $\psi_{n}(n=1,\cdots 5)$ with zero eigenvalues $\gamma_{n}=0$, contribute to the deviation function $\chi$ but remain unchanged with time. In the solution of Eq.(14), the second term, reflecting the contribution from collision invariants, can also be regarded as an integration constant. By comparing Eqs.(14) and (15), we can deduce that the traditional RTA simplifies the model by condensing the entire nonzero eigenvalue spectrum into a single representative eigenvalue, ${\gamma_{n}}\rightarrow\frac{1}{\tau_{R}}$. Given the exponentially decaying form, $1/\tau_{R}$ should be identified as the smallest eigenvalue, where the mode $\psi_{6}$ persists until the late stage of evolution. In other words, other modes with a larger eigenvalue, which should have been absent from an earlier time, extend their lifetime to the final stage in this approximation.